OpenStudy (anonymous):

The perimeter of a rectangle is 48 feet. If the length measures 16 feet, which is the width?

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 2(l + b) = 48 }\) \(\Large \color{purple}{\rightarrow 2(16 + b) = 48 }\) \(\Large \color{purple}{\rightarrow 32 + 2b = 48 }\) \(\Large \color{purple}{\rightarrow 2b = 16 }\) \(\Large \color{purple}{\rightarrow b = 8 }\) 8 is the width/breadth.

Parth (parthkohli):

Note that Perimeter of a rectangle = 2(length + breadth)

OpenStudy (anonymous):

thank u